### v1

@@ 2 When to expect a low-rank solution
We repeat the same experiment changing the shifts to $s_{j}=1+10^{-10}+i\omega_{j}10^{-k}$ so that $\mathbf{s}=1+10^{-10}$, $\epsilon=\mathcal{O}(10^{-k})$, and $\min_{j=1,\ldots,n}|\lambda_{j}+\mathbf{s}|=10^{-10}$. In Table 1 (right), we report the results. For this problem, Proposition 2.1 cannot guarantee the low-rank approximability of $X$. Indeed, we have $c(\epsilon,\mathbf{s},\Lambda)=\mathcal{O}(10^{10-k})$ which is large for the values of $k$ we tested. This is numerically confirmed by the large errors reported in Table 1 (right). On the other hand, $X$ can still be well-approximated by a rank-1 matrix as indicated by the small values of $\sigma_{2}(X)/\sigma_{1}(X)$. This suggests that further theoretical investigation is necessary to identify all the scenarios where we can have a numerically low-rank solution.
@@ 3 Minimal residual rational Krylov approach
Due to the diagonal pattern of $S$, we employ a projection method with only left projection for the solution of the Sylvester equation (5). In this framework, the sought after solution is of the form $X_{m}=V_{m}Y_{m}$ with the columns of $V_{m}=[v_{1},\ldots,v_{m}]\in\mathbb{C}^{n\times m}$ forming an orthonormal basis for a suitable subspace and $Y_{m}\in\mathbb{C}^{m\times\ell}$ computed by imposing convenient conditions. In this paper, we propose the columns of $V_{m}$ to span the rational Krylov subspace (7). The basis $V_{m}$ of (7) can be computed via the rational Arnoldi scheme [26] or the rational Lanczos method [20] in the case of Hermitian $A$. Notice that we always assume $A$ to be real, so that if $\xi_{j}\in\mathbb{R}$, for all $j$, $V_{m}$ is real as well. A real basis can also be constructed in the case of complex poles coming in conjugate pairs; see, e.g., [17]. We compute $Y_{m}=[y_{1},\ldots,y_{\ell}]\in\mathbb{C}^{m\times\ell}$ by imposing a minimal residual condition, namely
$Y_{m}$ $=\operatorname*{arg\,min}_{Y\in\mathbb{C}^{m\times\ell}}\|AV_{m}Y+V_{m}YS-b\mathbf{1}_{\ell}^{T}\|_{F}^{2}$
(9) $=\operatorname*{arg\,min}_{Y=[y_{1},\ldots,y_{\ell}]}\sum_{j=1}^{\ell}\|AV_{m}y_{j}+s_{j}V_{m}y_{j}-b\|_{2}^{2}.$
To this end, we exploit the following rational Arnoldi relation [26]
$AV_{m}$ $=V_{m}T_{m}+h_{m+1,m}\xi_{m+1}v_{m+1}e_{m}^{T}H_{m}^{-1}-h_{m+1,m}(I-V_{m}V_{m}^{T})Av_{m+1}e_{m}^{T}H_{m}^{-1}$
$=V_{m+1}\begin{bmatrix}T_{m}\\ h_{m+1,m}(\xi_{m+1}-t_{m+1,m+1})e_{m}^{T}H_{m}^{-1}\\ \end{bmatrix}+h_{m+1,m}(I-V_{m+1}V_{m+1}^{T})Av_{m+1}e_{m}^{T}H_{m}^{-1}$
$=V_{m+1}\begin{bmatrix}T_{m}\\ h_{m+1,m}(\xi_{m+1}-t_{m+1,m+1})e_{m}^{T}H_{m}^{-1}\\ \end{bmatrix}+h_{m+1,m}\alpha\widetilde{v}_{m+2}e_{m}^{T}H_{m}^{-1}$
(10) $=[V_{m+1},\widetilde{v}_{m+2}]\begin{bmatrix}T_{m}\\ h_{m+1,m}(\xi_{m+1}-t_{m+1,m+1})e_{m}^{T}H_{m}^{-1}\\ h_{m+1,m}\alpha e_{m}^{T}H_{m}^{-1}\\ \end{bmatrix},$
where $T_{m}=V_{m}^{T}AV_{m}$, $t_{m+1,m+1}=v_{m+1}^{T}Av_{m+1}$, $H_{m}\in\mathbb{R}^{m\times m}$ and $h_{m+1,m}\in\mathbb{R}$ are quantities related to the orthonormalization step performed for the basis construction, and
$\alpha\widetilde{v}_{m+2}=(I-V_{m+1}V_{m+1}^{T})Av_{m+1},\quad\alpha=\|(I-V_{m+1}V_{m+1}^{T})Av_{m+1}\|.$
Notice that $\widetilde{v}_{m+2}$ would be the $(m+2)$-th basis vector computed by selecting the pole $\xi_{m+2}=\infty$ and that this vector is orthogonal to $V_{m+1}$ by construction.
By using (3), we can write the residual matrix as follows
$R_{m}=$ $\,AV_{m}Y_{m}+V_{m}Y_{m}S-b\mathbf{1}_{\ell}^{T}$
(11) $=$ $\,[V_{m+1},\widetilde{v}_{m+2}]\left(\begin{bmatrix}T_{m}\\ h_{m+1,m}(\xi_{m+1}-t_{m+1,m+1})e_{m}^{T}H_{m}^{-1}\\ h_{m+1,m}\alpha e_{m}^{T}H_{m}^{-1}\end{bmatrix}Y_{m}+\begin{bmatrix}I_{m}\\ 0\\ \end{bmatrix}Y_{m}S-\beta e_{1}\mathbf{1}_{\ell}^{T}\right)$
@@
(12) $\|R_{m}\|_{F}^{2}=$ $\left\|\begin{bmatrix}T_{m}\\ h_{m+1,m}(\xi_{m+1}-t_{m+1,m+1})e_{m}^{T}H_{m}^{-1}\\ h_{m+1,m}\alpha e_{m}^{T}H_{m}^{-1}\end{bmatrix}Y_{m}+\begin{bmatrix}I_{m}\\ 0\\ \end{bmatrix}Y_{m}S-\beta e_{1}\mathbf{1}_{\ell}^{T}\right\|_{F}^{2},$
thanks to the orthogonality of $[V_{m+1},\widetilde{v}_{m+2}]$.
By using (12), the matrix $Y_{m}$ in (3) can thus be computed by solving the $(m+2)\times m$ least squares problem
(13) ${Y_{m}=\operatorname*{arg\,min}_{Y\in\mathbb{C}^{m\times m}}\left\|\begin{bmatrix}T_{m}\\ h_{m+1,m}(\xi_{m+1}-t_{m+1,m+1})e_{m}^{T}H_{m}^{-1}\\ h_{m+1,m}\alpha e_{m}^{T}H_{m}^{-1}\end{bmatrix}Y+\begin{bmatrix}I_{m}\\ 0\\ \end{bmatrix}YS-\beta e_{1}\mathbf{1}_{\ell}^{T}\right\|_{F}.}$
@@
(14) $y_{m}^{(j)}=\operatorname*{arg\,min}_{y\in\mathbb{C}^{m}}\left\|\begin{bmatrix}T_{m}+s_{j}I_{m}\\ h_{m+1,m}(\xi_{m+1}-t_{m+1,m+1})e_{m}^{T}H_{m}^{-1}\\ h_{m+1,m}\alpha e_{m}^{T}H_{m}^{-1}\end{bmatrix}y-\beta e_{1}\right\|.$
If
$[Q_{m}^{(j)},P_{m}^{(j)}]\begin{bmatrix}G_{m}^{(j)}\\ 0\\ \end{bmatrix}=\begin{bmatrix}T_{m}+s_{j}I_{m}\\ h_{m+1,m}(\xi_{m+1}-t_{m+1,m+1})e_{m}^{T}H_{m}^{-1}\\ h_{m+1,m}\alpha e_{m}^{T}H_{m}^{-1}\end{bmatrix},$
denotes the QR factorization of the coefficient matrix in (14) with $Q_{m}^{(j)}\in\mathbb{C}^{(m+2)\times m}$, $P_{m}^{(j)}\in\mathbb{C}^{(m+2)\times 2}$, and $G_{m}\in\mathbb{C}^{m\times m}$ upper triangular, then
$y_{m}^{(j)}=\beta G_{m}^{-1}((Q_{m}^{(j)})^{T}e_{1}),$
and the norm of the residual vector $r_{m}^{(j)}=(A+s_{j}I_{n})V_{m}y_{m}^{(j)}-b$ can be cheaply computed by
(15) $\|r_{m}^{(j)}\|=\left\|\begin{bmatrix}T_{m}+s_{j}I_{m}\\ h_{m+1,m}(\xi_{m+1}-t_{m+1,m+1})e_{m}^{T}H_{m}^{-1}\\ h_{m+1,m}\alpha e_{m}^{T}H_{m}^{-1}\end{bmatrix}y_{m}^{(j)}-\beta e_{1}\right\|=\beta\|(P_{m}^{(j)})^{T}e_{1}\|.$
It is important to stress that the vector $\widetilde{v}_{m+2}$ does not need to be explicitly computed. Indeed, for the definition of the projected least squares problems, we only need its norm $\alpha$, which can be easily calculated. Indeed,
$\alpha^{2}$ $\,=v_{m+1}^{T}A^{T}(I-V_{m+1}V_{m+1}^{T})^{2}Av_{m+1}=v_{m+1}^{T}A^{T}(I-V_{m+1}V_{m+1}^{T})Av_{m+1}$
(16) $\,=\|Av_{m+1}\|^{2}-\|V_{m+1}^{T}Av_{m+1}\|^{2},$
and the vectors $Av_{m+1}$ and $V_{m+1}^{T}Av_{m+1}$ can also be used in the following iteration to define $T_{m+1}=V_{m+1}^{T}AV_{m+1}$. Similarly, once the residual norm is sufficiently small, $\widetilde{v}_{m+2}$ is not required in the construction of the current approximate solution as $X_{m}=V_{m}Y_{m}$.
@@
1:Input: matrix $A\in\mathbb{R}^{n\times n}$, shifts $s_{i}\in\mathbb{C}$, right-hand side $b\in\mathbb{R}^{n}$, maximum number of iterations maxit, tolerance tol.
2:Output: $V_{m}\in\mathbb{C}^{n\times m}$, $Y_{m}\in\mathbb{C}^{m\times\ell}$ such that $\|(A+s_{i}I_{n})V_{m}Y_{m}e_{i}-b\|\leq\|b\|\cdot\texttt{tol}$
3:Set $V_{1}=v_{1}=b/\|b\|$ and select first pole $\xi_{2}$
4:Set $\Sigma=\{1,\ldots,\ell\}$ and $\Sigma_{C}=\emptyset$
@@
6:     Solve $(A+\xi_{m+1}I_{n})w=v_{m}$
7:     Compute $v_{m+1}$ by orthonormalizing $w$ wrt $V_{m}$
8:     Set $V_{m+1}=[V_{m},v_{m+1}]$ and compute $T_{m+1}=V_{m+1}^{T}AV_{m+1}$ and $\alpha$ as in (3)
9:     for $j\in\Sigma\setminus\Sigma_{C}$ do
10:         Compute $y_{m}^{(j)}$ as in (14)
11:         Compute the residual norm $\|r_{m}^{(j)}\|$ as in (15)
12:         if $\|r_{m}^{(j)}\|<\|b\|\cdot\mathtt{tol}$ then
13:              Set $\Sigma_{C}=\Sigma_{C}\cup\{j\}$ and $Y_{m}e_{j}=y_{m}^{(j)}$
14:         end if
15:         if $\Sigma=\emptyset$ then
16:              Go to line 21
17:         end if
18:     end for
19:     Compute next pole $\xi_{m+2}$
20:end for
21:Return $V_{m}$ and $Y_{m}$ by possibly padding the latter with zeros
@@
We remind the reader that the orthonormalization step in line 7 yields the matrix $H_{m}$ and the scalar $h_{m+1,m}$ utilized in the definition of the projected problem (14). Moreover, notice that the nesting feature of rational Krylov subspaces, along with the appealing residual minimization property of our routine, guarantees that whenever the residual norm related to the $j$-th linear system in (1), namely $\|r_{m}^{(j)}\|$, is smaller than the prescribed accuracy target, it remains so also when the subspace is expanded as $\|r_{m+k}^{(j)}\|\leq\|r_{m}^{(j)}\|$ for any $k\geq 0$. This can be exploited in computation. In particular, if at iteration $m$ we detect a sufficiently small residual norm for the $j$-th linear system, we stop solving the projected problem related to that linear system in the following iterations. By doing so, the number of small-dimensional least squares problems (14) can potentially decrease along the rational Krylov iterations. On the other hand, this strategy requires careful handling of the matrix $Y_{m}$. In particular, if at the $m$-th iteration $\|r_{m}^{(j)}\|$ is sufficiently small, then we set $Y_{m}e_{j}=y_{m}^{(j)}$. However, it is very unlikely that at this point all the linear systems have converged, namely $\|r_{m}^{(i)}\|$ is sufficiently small for all $i=1,\ldots,\ell$. We thus proceed with expanding the space but we no longer solve the least squares problem related to the $j$-th linear system. Therefore, for any $k\geq 0$ we set $Y_{m+k}e_{j}=[y_{m}^{(j)};0]\in\mathbb{C}^{m+k}$ where, with abuse of notation, 00 denotes a $k$-dimensional vector of all zeros.
@@
Algorithm 1 is given by assuming that all the shifts $s_{i}$ are available. However, working in a framework where new shifts are provided every now and then does not lead to any particular difficulties. Indeed, the construction of the rational Krylov subspace is independent of the shifts $s_{i}$’s. Whenever a new shift comes at hand, one can simply assemble the related projected least squares problem (14) with the rational Krylov quantities already available.

### v2

@@ 2 When to expect a low-rank solution
We repeat the same experiment changing the shifts to $s_{j}=1+10^{-10}+i\omega_{j}10^{-k}$ so that $\mathbf{s}=1+10^{-10}$, $\epsilon=\mathcal{O}(10^{-k})$, and $\min_{j=1,\ldots,n}|\lambda_{j}+\mathbf{s}|=10^{-10}$. In Table 1 (right), we report the results. For this problem, Proposition 2.1 cannot guarantee the low-rank approximability of $X$. Indeed, we have $c(\epsilon,\mathbf{s},\Lambda)=\mathcal{O}(10^{10-k})$ which is large for the values of $k$ we tested. This is numerically confirmed by the large errors reported in Table 1 (right). Nevertheless, $X$ can still be well-approximated by a rank-1 matrix as indicated by the small values of $\sigma_{2}(X)/\sigma_{1}(X)$. This suggests that further theoretical investigation is necessary to identify all the scenarios where we can have a numerically low-rank solution.
Remark 2.4.
We mention here that we employ the concept of centroids only for theoretical purposes, to depict scenarios where we do expect the singular values of the exact solution to rapidly decay to zero. While centroids might be employed as poles in the rational Krylov basis construction, this strategy is likely not effective in settings where the results from Proposition 2.1 do not apply, cf. Example 2.3. Therefore, in section 3.1 we propose a general, greedy pole-selection strategy where the poles $\xi_{i}$’s in (7) are computed on the fly during the iterative process, as it is customary in rational Krylov subspace methods.
Remark 2.5.
In [KressnerTobler11] the authors report some results on the numerical linear dependency of solutions to parametrized linear systems of the form $A(\alpha)x(\alpha)=b(\alpha)$ where $\alpha\in[-1,1]$, and in [Rec1, Section 6.6] recycling Krylov methods have been tested in this scenario. In particular, in [KressnerTobler11, Theorem 2.4] it has been shown how the decay of the singular values of the matrix $x(\alpha_{1}),\ldots,x(\alpha_{p})]$ is bounded from above by $\max_{\alpha\in\mathcal{E}_{\rho}}\|A(\alpha)^{-1}\|_{2}$ where $\mathcal{E}_{\rho}\subset\mathbb{C}$ denotes the open elliptic disc with foci $\pm 1$ and the sum of the half axes equal to $\rho$. This result can be certainly applied to our context as well. By considering the technical assumption that $s_{i}\in[-1,1]$ for all $i=1,\ldots,\ell$, [KressnerTobler11, Theorem 2.4] says that the decay of the singular values of the solution $X$ to (5) is mainly driven by $\max_{s\in\mathcal{E}_{\rho}}\|(A+sI)^{-1}\|_{2}\geq\max_{i=1,\ldots,\ell}\|(A+s_{i}I)^{-1}\|_{2}$ underlying the importance of having well-conditioned problems in (1) but overlooking, once again, the possible clustering of the shifts $s_{i}$’s.
By building upon [KressnerTobler11], it might be interesting to study how (and if) a possible clustering of the parameters $\alpha$ impacts the decay of the singular values of $x(\alpha_{1}),\ldots,x(\alpha_{p})]$ thus extending our findings to general parameter-dependent problems $A(\alpha)x(\alpha)=b(\alpha)$. This fascinating topic is left to be studied elsewhere.
@@ 3 Minimal residual rational Krylov approach
Due to the diagonal pattern of $S$, we employ a projection method with only left projection for the solution of the Sylvester equation (5). In this framework, the sought after solution is of the form $X_{m}=V_{m+1}\underline{K}_{m}Y_{m}$ with the columns of $V_{m+1}=[v_{1},\ldots,v_{m+1}]\in\mathbb{C}^{n\times(m+1)}$ forming an orthonormal basis for a suitable subspace and $Y_{m}\in\mathbb{C}^{m\times\ell}$ computed by imposing convenient conditions. In this paper, we propose the columns of $V_{m+1}$ to span the rational Krylov subspace (7). The basis $V_{m+1}$ of (7) can be computed via the rational Arnoldi scheme [Ruhe94] or the rational Lanczos method [ratLanczos] in the case of Hermitian $A$. Notice that if $A$ is real, and $\xi_{j}\in\mathbb{R}$, for all $j$, $V_{m+1}$ is real as well. A real basis can also be constructed if $A$ is real and the complex poles come in conjugate pairs; see, e.g., [realbasis]. We compute $Y_{m}=[y_{1},\ldots,y_{\ell}]\in\mathbb{C}^{m\times\ell}$ by imposing a minimal residual condition, namely
$Y_{m}$ $=\operatorname*{arg\,min}_{Y\in\mathbb{C}^{m\times\ell}}\|AV_{m+1}\underline{K}_{m}Y+V_{m+1}\underline{K}_{m}YS-b\mathbf{1}_{\ell}^{T}\|_{F}^{2}$
(12) $=\operatorname*{arg\,min}_{Y=[y_{1},\ldots,y_{\ell}]}\sum_{j=1}^{\ell}\|AV_{m+1}\underline{K}_{m}y_{j}+s_{j}V_{m+1}\underline{K}_{m}y_{j}-b\|_{2}^{2}.$
Using the rational Arnoldi relation (8) we can write the residual matrix as follows
$R_{m}=$ $\,AV_{m+1}\underline{K}_{m}Y_{m}+V_{m+1}\underline{K}_{m}Y_{m}S-b\mathbf{1}_{\ell}^{T}$
(13) $=$ $\,V_{m+1}\left(\underline{H}_{m}Y_{m}+\underline{K}_{m}Y_{m}S-\beta e_{1}\mathbf{1}_{\ell}^{T}\right),$
@@
(14) $\|R_{m}\|_{F}^{2}=$ $\left\|\underline{H}_{m}Y_{m}+\underline{K}_{m}Y_{m}S-\beta e_{1}\mathbf{1}_{\ell}^{T}\right\|_{F}^{2},$
since $V_{m+1}$ is unitary. Therefore, the matrix $Y_{m}$ in (3) can be computed by solving the $(m+1)\times m$ least squares problem
(15) ${Y_{m}=\operatorname*{arg\,min}_{Y\in\mathbb{C}^{m\times m}}\left\|\underline{H}_{m}Y+\underline{K}_{m}YS-\beta e_{1}\mathbf{1}_{\ell}^{T}\right\|_{F}.}$
@@
(16) $y_{m}^{(j)}=\operatorname*{arg\,min}_{y\in\mathbb{C}^{m}}\left\|(\underline{H}_{m}+s_{j}\underline{K}_{m})y-\beta e_{1}\right\|.$
Using the QR decomposition of the coefficient matrix in the least-squares problem (16), we have
$Q_{m}^{(j)},P_{m}^{(j)}]\begin{bmatrix}G_{m}^{(j)}\\ 0\\ \end{bmatrix}=\underline{H}_{m}+s_{j}\underline{K}_{m},$
where $Q_{m}^{(j)}\in\mathbb{C}^{(m+1)\times m}$, $P_{m}^{(j)}\in\mathbb{C}^{(m+1)}$, and $G_{m}^{(j)}\in\mathbb{C}^{m\times m}$ upper triangular. Therefore,
$y_{m}^{(j)}=\beta(G_{m}^{(j)})^{-1}((Q_{m}^{(j)})^{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}*}e_{1}),$
and the norm of the residual vector $r_{m}^{(j)}=(A+s_{j}I_{n})V_{m}y_{m}^{(j)}-b$ can be cheaply computed as
(17) $\|r_{m}^{(j)}\|=\|(\underline{H}_{m}+s_{j}\underline{K}_{m})y_{m}^{(j)}-\beta e_{1}\|=\beta\|(P_{m}^{(j)})^{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}*}e_{1}\|.$
@@
1:Input: matrix $A\in{\mathbb{C}}^{n\times n}$, shifts $s_{i}\in\mathbb{C}$, right-hand side $b\in\mathbb{R}^{n}$, maximum number of iterations maxit, tolerance tol.
2:Output: $V_{m+1}\in\mathbb{C}^{n\times(m+1)}$, $Z_{m}\in\mathbb{C}^{(m+1)\times\ell}$ such that $\|(A+s_{i}I_{n})V_{m+1}Z_{m}e_{i}-b\|\leq\|b\|\cdot\texttt{tol}$
3:Set $V_{1}=v_{1}=b/\|b\|$
4:Set $\Sigma=\{1,\ldots,\ell\}$, $\Sigma_{C}=\emptyset$, and randomly select an initial pole $\xi_{1}$.
@@
6:  Solve $(A+\xi_{m}I_{n})w=v_{m}$
7:  Set up $\underline{K}_{m}$ and $\underline{H}_{m}$ so that $AV_{m+1}\underline{K}_{m}=V_{m+1}\underline{H}_{m}$
8:  for $j\in\Sigma\setminus\Sigma_{C}$ do
9:   Compute $y_{m}^{(j)}$ as in (16) and set $z_{m}^{(j)}=\underline{K}_{m}y_{m}^{(j)}$
10:   Compute the residual norm $\|r_{m}^{(j)}\|$ as in (17)
11:   if $\|r_{m}^{(j)}\|<\|b\|\cdot\mathtt{tol}$ then
12:     Set $\Sigma_{C}=\Sigma_{C}\cup\{j\}$ and $Z_{m}e_{j}=z_{m}^{(j)}$
13:   end if
14:   if $\Sigma\setminus\Sigma_{C}=\emptyset$ then
15:     Go to line 20
16:   end if
17:  end for
18:  Compute next pole $\xi_{m+1}$
19:end for
20:Return $V_{m+1}$ and $Z_{m}$ by possibly padding the latter with zeros
@@
Note that since, $\text{Range}(V_{i+1}\underline{K}_{i})\subseteq\text{Range}(V_{j+1}\underline{K}_{j})$ for $i\leq j$, the appealing residual minimization property of our routine guarantees that whenever the residual norm related to the $j$-th linear system in (1), namely $\|r_{m}^{(j)}\|$, is smaller than the prescribed accuracy target, it remains so also when the subspace is expanded as $\|r_{m+k}^{(j)}\|\leq\|r_{m}^{(j)}\|$ for any $k\geq 0$.
@@
The monotone decrease of the residual norm can be exploited in computation. In particular, if at iteration $m$ we detect a sufficiently small residual norm for the $j$-th linear system, we stop solving the projected problem related to that linear system in the following iterations. By doing so, the number of small-dimensional least squares problems (16) can potentially decrease along the rational Krylov iterations. Note that this strategy requires careful handling of the matrix $Y_{m}$. In particular, if at the $m$-th iteration $\|r_{m}^{(j)}\|$ is sufficiently small, then we set $Y_{m}e_{j}=y_{m}^{(j)}$. However, it is very unlikely that at this point all the linear systems have converged, namely $\|r_{m}^{(i)}\|$ is sufficiently small for all $i=1,\ldots,\ell$. We thus proceed with expanding the space but we no longer solve the least squares problem related to the $j$-th linear system. Therefore, for any $k\geq 0$ we set $Y_{m+k}e_{j}=[y_{m}^{(j)};0]\in\mathbb{C}^{m+k}$ where, with abuse of notation, 0 denotes a $k$-dimensional vector of all zeros.
Remark 3.2.
For the sake of simplicity, Algorithm 1 is given by assuming that all the shifts $s_{i}$ are available. However, working in a framework where new shifts are provided every now and then does not lead to any particular difficulties. Indeed, the construction of the rational Krylov subspace is independent of the shifts $s_{i}$’s. Whenever a new shift comes at hand, one can simply assemble the related projected least squares problem (16) with the rational Krylov quantities already available. This means that we can fully capitalize on the computational efforts done so far also for the solution of linear systems with newly acquired shifts and the solution process does not need to start from scratch all over again.
Linear systems solution $\mathcal{O}(m\cdot\text{nnz}(A))$
Orthogonalization $\mathcal{O}(nm^{2})$
Solution of (16) $\mathcal{O}(\ell m^{4})$
Table 2: Main asymptotic costs (in flops) of performing $m$ iterations of Algorithm 1.
We conclude this section by reporting in Table 2 the main asymptotic costs of Algorithm 1, after $m$ iterations. We remind the reader that, at iteration $k$, the cost per iteration of the orthogonalization step and the solution of the projected problems (16) amounts to $\mathcal{O}(nk)$ and $\mathcal{O}(\ell k^{3})$ flops, respectively, so that after $m$ iterations we end up with the values reported in Table 2. These values highlight the importance of constructing a very meaningful approximation space, able to converge in a few iterations. Otherwise, Algorithm 1 would not be competitive with the naive solution of the $\ell$ linear systems in (1), whose cost is $\mathcal{O}(\ell\cdot\text{nnz}(A))$ flops. Similarly, the strategy mentioned in Remark 3.1 can significantly help reducing the cost of the solution of the projected problems as not $\ell$ least squares problems need to be solved at each iteration. The cost reported in Table 2 is the worst-case scenario, where we have to solve all the $\ell$ projected problems in (16) until the $m$-th iteration.
Remark 3.3.
Note that the QR decomposition of $\underline{H}_{m}+s_{j}\underline{K}_{m}$ for each $j=1,\ldots,\ell$, can be updated at each iteration by using a Givens rotations approach similar to how the latter is employed in GMRES [Schultz1986] and the residual norm computation can be performed almost for free reducing the cost of (16) to $\mathcal{O}(\ell m^{2})$ when the solution $Y_{m}$ is only computed once the residual is below the required threshold. However, this reduction in cost comes with the extra memory requirement to store additional $\ell$ matrices of size $(m+1)\times m$, the $R$ factors; $\ell$ vectors of length $m$, the residuals; and $\ell$ sets of Givens rotations each having $m$ rotations.